Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2021-05-02

Observers: Emmy Paraskeva, Miika Pursiainen
Support: Daniele Malesani, Giorgos Leloudas
TAT: Lluís Galbany, Roberta Carini, Shubham Srivastav

EFOSC Calibrations: 
Bias, dome_flats, arcs.


Weather:
Start: Some thin clouds, seeing 1.1", wind from north 10 m/s, humidity ~30%
Middle:  Strong wind: 16 m/s, seeing 2"
, humidity ~30%, Clouds getting thicker
End: Strong wind up to 25 m/s, humidity 40%, thin clouds.

Comments: Issue with telescope rotor at UT=00:25

EFOSC


 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
23:15
L745aSTD



1.1
1.1Image analysis performed on the STD
23:35
SN2019qemFollow-up
g13
1"
1x2700s
1.5
1.1
00:25
Issue with Telescope rotor!

  



Telescope rotor got stuck when moving to new target
00:50
SN2021bnw
Follow-upg13
1"
1x2700s 1.4
1.4
Wind increased significantly during observation possibly affecting the exposure.
01:45
TELESCOPE CLOSED






Telescope was forced to close due to high wind









Comments